OpenStudy (nightmarenight):

distributive property -(x + 3)

OpenStudy (nincompoop):

so the negative or minus sign indicate that there is \(-1 \times (x+3) \)

OpenStudy (nincompoop):

and distributive property has the following general pattern \(a(b\color{red}{+} c) = (a\times b) \color{red}{+}(a \times c) \) an example would be \(-2(3-4) = (-2 \times 3 ) \color{red}{-}(-2 \times -4)\)